How would you enter the inequality 2x - 3y >= 0 into the calculator to graph it?

Answer by stanbon(48516)

How would you enter the inequality 2x - 3y >= 0 into the calculator to graph it?
Rewrite the inequality as:
2x>=3y
y<=(2/3)x
Enter y=(2/3)x into the calculator to get the line y=(2/3)x
The "<" means to take all the points below the line.
Cheers,
Stan H.